# DiffOG: Differentiable Policy Trajectory Optimization with Generalizability

### v3

@@ V-A Benchmark with Baselines Across Tasks
4. For tasks where trajectory constraints and smoothness are crucial, such as move the stack in Table III, the base policy often fails due to the jerkiness of the action trajectories, resulting in a high failure rate. In contrast, DiffOG significantly improves trajectory quality, leading to better policy performance.
@@
V-B Ablation Study on Adjustability of Trajectory
@@
V-C Ablation Study on Static $\mathbf{Q}$, Matrix Learning, and Transformer
@@
We selected the transport task as the test task. This is a dual-arm, long-horizon manipulation task, which poses challenges for trajectory optimization in terms of generalizing to trajectories of varying forms within long horizon manipulation and high-dimensional action space. The experimental results are shown in Table V. From the table, we observe that DiffOG refine training achieves the best performance.
The explanation behind this is that for static $\mathbf{Q}$ and matrix-learning-based $\mathbf{Q}$, trajectory optimization cannot effectively adjust to different trajectory inputs. This highlights the flexibility and generalization capability of DiffOG, thanks to the transformer-based embedding design.
@@
TABLE V: Success rate of different $\mathbf{Q}$ designs on transport task.
V-D Comparison with a Constrained Visuomotor Policy
We conducted comparison experiments between DiffOG and constrained visuomotor policy Leto [32] on 7 tasks. The results are shown in Table VI. The results in the table compare DiffOG and Leto under the same constraint bounds.
Based on the results in Table VI, we can see that the performance of policies enhanced by DiffOG is entirely superior to Leto policies. This demonstrates the generalization capability of DiffOG across diverse tasks.
@@ V-D Comparison with a Constrained Visuomotor Policy
TABLE VI: Success rate of benchmarking DiffOG with Leto [32].
VI Limitations and Future Work
@@
VII Conclusion
@@
Lift 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 $2\times 3\times 84\times 84$
Can 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 $2\times 3\times 84\times 84$
Square 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 $2\times 3\times 84\times 84$
Tool Hang 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 $2\times 3\times 240\times 240$
Transport 0.1 0.1 200 12-DOF Vel. + 2-DOF Grasping 0.05 s 4 $4\times 3\times 84\times 84$
Push-T 0.05 [12.5,12.2] Pixels 90 2-DOF Position 0.1 s 1 $1\times 3\times 96\times 96$
Pick Place Wall 0.6 [0.53, 0.92, 3.75] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Shelf Place 0.6 [2.32, 1.47, 4.13] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Disassemble 0.6 [0.65, 1.10, 0.61] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Stick Push 0.6 [1.03, 0.15, 0.68] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Stick Pull 0.6 [4.46, 1.44, 2.49] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Arrange Desk 0.1 [0.070, 0.071, 0.102, 0.030, 0.056, 0.148, 0.078, 0.089, 0.105, 0.039, 0.069, 0.120] rad 100 12-DOF Angle + 2-DOF Grasping 0.25 s 4 $3\times 3\times 240\times 320$
Move the Stack 0.05 [0.049, 0.051, 0.057, 0.045, 0.038, 0.069] rad 100 6-DOF Angle + 1-DOF Grasping $\frac{1}{6}$ s 4 $2\times 3\times 240\times 320$
@@
TABLE VII: Summary of hyperparameters for trajectory optimization and dataset. $\alpha$ is the smooth weight, $\frac{1}{\Delta t}$ represents the frequency of policy inference, and “Demos” indicates the number of demonstration episodes. “High Dim. Obs” refers to the dimensionality of high-dimensional observations. For instance, $2\times 3\times 84\times 84$ corresponds to RGB observations from two cameras, each with a shape of $3\times 84\times 84$, while $3\times 512$ represents point cloud data with a shape of $3\times 512$. “Bound” refers to $d_{\text{max}}\Delta t$ in equation (7). “Unnormalized Bound” refers to constraint bounds unnormalized to dataset scale.
Appendix
VII-A Hyperparameters and Dataset
@@
The summary of hyperparameters for trajectory optimization and dataset as shown in Table VII. In the experiments of this paper, we consistently adopt symmetric constraints which is $d_{\text{max}}=-d_{\text{min}}$. Since the actions require normalization to [-1,1] during training [5], all constraints are applied on the normalized action scale for DiffOG. For simplicity, when applying constraints on the normalized action scale, the same value is used for each action dimension. Therefore, the “Bound” column contains only a single value for each task. Since the distribution of action values differs across task datasets, the bounds for each action dimension may vary after unnormalization. For Robomimic tasks [12], the action values in the dataset are already normalized to the range of $[-1,1]$. Therefore, the bound and the unnormalized bound are identical.
@@
Since constraint clipping and penalty-based trajectory optimization do not require training, constraints can be directly applied to unnormalized actions. These constraints are equivalent to those imposed by DiffOG but are applied directly as unnormalized constraints. The parameters for penalty-based trajectory optimization are listed in Table VIII. As mentioned earlier, the actions provided by the Robomimic dataset are already normalized to the range of $[-1,1]$. The constraint bound we impose in this case is $0.1$. Based on this, the violation tolerance for penalty-based optimization is set to a relatively small value of $0.004$. For other tasks, the constraints are applied to raw action values, and the violation tolerances are determined based on the scale of unnormalized bounds provided in Table VII.
Finally, we provide the hyperparameters of the transformer used in DiffOG, as shown in Table IX.
@@
TABLE VIII: Parameters of penalty-based trajectory optimization.
@@
TABLE IX: Hyperparameters of transformer encoder.
VII-B Computation of Trajectory Metrics

### v4

@@ IV-A Benchmark with Baselines Across Tasks
4. For tasks where trajectory constraints and smoothness are crucial, such as move the stack in Table III, the base policy often fails due to the jerkiness of the action trajectories, resulting in a high failure rate. The jerky behavior of base policy mainly arises from two factors: 1) lack of explicit trajectory constraints, and 2) absence of smoothness priors in the policy output. These are common issues in end-to-end neural policies, especially for manipulation tasks requiring smooth, precise, and constraint-aware motions. In contrast, DiffOG significantly improves trajectory quality, leading to better policy performance.
@@
IV-B Ablation Study on Adjustability of Trajectory
@@
While the smoothing weight $\alpha$ and constraint bounds are manually selected, their selection can be guided by interpretable trends revealed in Fig. 9. Rather than hindering practical deployment, their interpretability, robustness, and task-specific tunability make DiffOG a flexible and practical framework for real-world use.
IV-C Ablation Study on Static $\mathbf{Q}$, Matrix Learning, and Transformer
@@
We selected the tool hang and transport as the test tasks. These are highly challenging long-horizon manipulation tasks, which make trajectory optimization difficult due to the need to generalize across diverse trajectory patterns over extended time horizons while maintaining precise control. The experimental results are shown in Table V. From the table, we observe that DiffOG achieves the best performance for both tasks.
The explanation behind this is that for static $\mathbf{Q}$ and matrix-learning-based $\mathbf{Q}$, trajectory optimization cannot effectively adjust to different trajectory inputs. This highlights the flexibility and generalization capability of DiffOG through transformer-based embeddings.
@@
Moreover, as shown in Table VI, the computational costs of solving a static $\mathbf{Q}$, optimizing with a matrix-based representation (matrix learning), and solving the transformer-based optimization (DiffOG) are comparable. This suggests that the additional inference time compared to base policy primarily stems from solving the optimization problem, rather than from the transformer forward pass.
@@
TABLE V: Success rate of different $\mathbf{Q}$ designs on tool hang and transport task.
Base Policy: Diffusion Policy DiffOG + Base Static $\mathbf{Q}$ + Base Matrix Learning + Base
Batch Inference Time 0.13 s 0.28 s 0.28 s 0.27 s
TABLE VI: We report the batched average inference time across three simulated tasks: lift, can, and square. Each measurement is based on a batch size of 25 and averaged over 50 inference runs. All methods are evaluated using 100 DDPM steps [9].
[…]
Figure 10: Diagonality score of the learned cost matrix $\mathbf{Q}$ throughout the tool hang process. $\mathbf{Q}$ becomes more diagonal during fine-manipulation phases (B: grasping, D: insertion, F: grasping, H: hanging), and less diagonal during fast transitions, revealing how DiffOG adapts optimization structure to task dynamics.
[…]
IV-D Comparison with a Constrained Visuomotor Policy
We conducted comparison experiments between DiffOG and constrained visuomotor policy Leto [33] on 7 tasks. The results are shown in Table VII. The results in the table compare DiffOG and Leto under the same constraint bounds.
Based on the results in Table VII, we can see that the performance of policies enhanced by DiffOG is superior to Leto policies. This demonstrates the generalization capability of DiffOG across diverse tasks.
@@ IV-D Comparison with a Constrained Visuomotor Policy
TABLE VII: Success rate of benchmarking DiffOG with Leto [33].
IV-E Inference Time-varying Constraint Experiment
Varying Constraints 0.05–0.3 Static Constraints 0.1
P1 Acc. P2 Acc. P3 Acc. P4 Acc. SR Acc. SR
Can 0.05/0.013 0.08/0.022 0.13/0.034 0.22/0.043 0.98 0.09/0.026 0.98
Square 0.05/0.013 0.07/0.018 0.09/0.023 0.17/0.028 0.83 0.09/0.020 0.87
TABLE VIII: Inference time-varying constraint experiment (Phases P1–P4 with bounds 0.05, 0.1, 0.2, and 0.3, respectively). The policy is trained on a fixed normalized constraint bound of 0.1. During inference, tighter bounds result in smoother motion, reflected by lower acceleration standard deviation and capped maximum acceleration, while preserving a success rate comparable to that of the policy using static constraints 0.1. Acc. represents acceleration and SR represents success rate. Accelerations here refer to maximum/standard deviation.
[…]
Figure 11: Task phase illustration in the inference time-varying constraint experiment.
In this experiment, the policy is trained on a fixed normalized constraint bound of 0.1 for the can and square tasks. However, during inference, we apply time-varying bounds: 0–40 steps use 0.05 (Phase 1), 40–80 use 0.1 (P2), 80–120 use 0.2 (P3), and 120 to task completion use 0.3 (P4). The task phases used for applying varying constraints during inference are illustrated in Fig. 11. Note that the phases are determined using uniform intervals of 40 steps, without any task-specific selection, as our goal is to evaluate the low-level capability of handling varying constraints. As shown in Table VIII, despite being trained only with a static constraint, DiffOG can generalize zero-shot to time-varying constraints at inference time due to its interpretability and generalizability.
IV-F Residual Policy Baseline with Smoothness Penalty
[…]
Figure 12: Residual policy training with varying smoothness penalty weight $\alpha$.
In this section, we design a residual policy training baseline that allows for smoothness regularization. Specifically, given the base action $\mathbf{a}_{t}$, we train a lightweight MLP to predict a residual action $\Delta\mathbf{a}_{t}=\mathrm{MLP}(\mathbf{a}_{t})$, and define the final action as $\mathbf{y}_{t}=\mathbf{a}_{t}+\Delta\mathbf{a}_{t}$. The training loss is
$\mathcal{L}_{p}=\mathbb{E}_{\mathbf{a}_{t}\in\mathcal{D}}\left[\|\mathbf{y}_{t}-\mathbf{a}_{t}\|^{2}+\frac{\alpha}{2}\left(\frac{d\hat{\mathbf{c}}_{t}}{dt}\right)^{\!\!\mathrm{T}}\left(\frac{d\hat{\mathbf{c}}_{t}}{dt}\right)\right],$
where the second term is identical to the smoothness penalty used in DiffOG and encourages low velocity/acceleration over the resulting trajectory. The smoother the motion, the smaller this term becomes. Please refer to equations (5) and (6) for implementation details. We conduct extensive ablation experiments by varying $\alpha$ across a wide range (from 0.01 to 16) and evaluate on three tasks: lift, can, and square. As shown in Fig. 12, we observe the following: 1) when $\alpha$ is small (e.g., 0.01), the smoothness penalty is relatively weak, and the residual policy maintains comparable success rates to both DiffOG and the base policy; 2) under the same setting, the residual reduces max and std acceleration over the base policy, indicating that smoother actions are indeed being generated; 3) however, the improvement in smoothness still falls significantly short of DiffOG, with noticeably higher max and std acceleration of residual policy baseline; 4) as $\alpha$ increases, task success rate consistently drops across all three tasks; 5) although larger $\alpha$ generally leads to smoother motions, the gains quickly saturate, and the resulting trajectories remain less smooth than those produced by DiffOG, even when the success rate degrades substantially (e.g., $\alpha>4$); and 6) in some cases, further increasing $\alpha$ even slightly increase the max or std acceleration. These results indicate that DiffOG is a more effective solution, achieving substantial trajectory improvement without sacrificing policy performance.
In addition to its empirical advantages, DiffOG offers a fundamental distinction from the residual policy with smoothness regularization. The residual policy cannot guarantee hard constraint satisfaction at inference time and lacks a structured, interpretable optimization framework. In contrast, DiffOG leverages differentiable trajectory optimization to explicitly enforce hard constraints and systematically refine trajectories. The structure of DiffOG leads to better alignment with task requirements and stronger overall performance, making DiffOG a more principled and generalizable method.
V Limitations and Future Work
@@
Moreover, our method has demonstrated strong empirical performance under widely adopted action representations in imitation learning, such as joint angles [36, 29] and 6D end-effector velocities [13, 27, 37]. Beyond these settings, we identify several promising directions for future work: 1) Certain action spaces that are less common in current imitation learning benchmarks, such as joint torques, also conform to the formulations of DiffOG. While not yet widely used, joint torques hold strong potential for force-based imitation learning, as they provide direct control over physical interactions. Extending our method to torque-based action spaces could enable more contact-rich manipulation. 2) Future extensions of the framework may address more complex representations, such as absolute $SO(3)$ rotations. In such cases, the assumption of simple discrete-time derivatives may no longer hold, requiring more expressive formulations of the underlying optimization problem.
In addition, a natural extension of this work is to incorporate collision avoidance constraints into the DiffOG framework. This would require integrating environment and geometry information to formulate differentiable constraints that ensure safe interaction with surrounding objects. While this introduces additional complexity in constraint modeling, it would significantly broaden the applicability of DiffOG to tasks involving cluttered or dynamic environments.
Lastly, DiffOG demonstrates the ability to perform inference with time-varying constraints, even when trained only with static ones. If a high-level semantic reasoning module is available to generate time-varying constraints, it could be seamlessly integrated with the DiffOG formulation to form a hierarchical structure. This would enable more phase-specific optimization tailored to different stages of a task, representing a promising direction for future work.
VI Conclusion
@@
Appendix
VI-A Baselines
where $\mathbf{S}$ is for selecting the action dimensions to be considered for trajectory optimization. Similarly, we define an output variable
VI-A1 Greedy Constraint Clipping
The optimized action at time $t+k$ is then updated as
$\hat{c}_{t+k}^{\star}=\hat{c}_{t+k-1}^{\star}+\Delta c_{t+k}^{\text{clip}}.$
VI-A2 Penalty-Based Trajectory Optimization
$\nabla=\mathbf{v}_{\text{max}}-\mathbf{v}_{\text{min}}.$
${\mathbf{\hat{c}}}_{t}\leftarrow{\mathbf{\hat{c}}}_{t}-\eta\nabla,$
VI-B Action Space Assumption
Assumption 1.
Another widely adopted action space is $[v_{x},\,v_{y},\,v_{z},\,\dot{r}_{x},\,\dot{r}_{y},\,\dot{r}_{z}]^{\mathrm{T}}\in\mathbb{R}^{6},$ where $\mathbf{v}_{xyz}=[v_{x},v_{y},v_{z}]^{\mathrm{T}}$ is translational velocity, and $\dot{\mathbf{r}}=[\dot{r}_{x},\dot{r}_{y},\dot{r}_{z}]^{\mathrm{T}}$ is the derivative of the axis-angle rotation vector [13, 27, 37, 33]. Its time derivative, $[\dot{v}_{x},\,\dot{v}_{y},\,\dot{v}_{z},\,\ddot{r}_{x},\,\ddot{r}_{y},\,\ddot{r}_{z}]^{\!\mathrm{T}},$ includes linear accelerations and axis-angle second derivatives, which differ from the true angular acceleration $\dot{\bm{\omega}}$.
Although $\ddot{\mathbf{r}}$ does not directly represent $\dot{\bm{\omega}}$, constraining $\ddot{\mathbf{r}}$ effectively regulates $\dot{\bm{\omega}}$. The angular acceleration is related by $\dot{\bm{\omega}}=\mathbf{J}(\mathbf{r})\ddot{\mathbf{r}}+\dot{\mathbf{J}}(\mathbf{r})\dot{\mathbf{r}}$, where $\mathbf{J}(\mathbf{r})$ is the Jacobian matrix mapping axis-angle velocity to angular velocity. In practice, $\mathbf{J}(\mathbf{r})$ typically remains bounded, as evidenced by empirical successes of such action space in imitation learning frameworks [13, 27, 37], which achieve stable rotations even without explicit constraints. Different from these prior work, DiffOG explicitly and effectively constrains rotational motion, resulting in smoother and more controlled behaviors. Extensive experimental results demonstrate the effectiveness of DiffOG.
VI-C Task Descriptions and Experimental Setup
For the end-effector 6-DOF velocity action space, we include five tasks: lift, can, square, tool hang, and transport, all sourced from Robomimic benchmark [13]. For the end-effector linear position action space, we include six tasks: push-T from [5], along with pick place wall, shelf place, disassemble, stick push, and stick pull from Meta-World [34]. Notably, we selected all the “very hard” tasks identified in Meta-World as marked in [35]. Lastly, for the joint angle action space, we include two real-world tasks, arrange desk and move the stack, with data collection and policy experiments conducted using ALOHA [36].
@@ VI-C Task Descriptions and Experimental Setup
Lift 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 2×3×84×842$\times$ 3$\times$ 84$\times$ 842 × 3 × 84 × 84
Can 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 2×3×84×842$\times$ 3$\times$ 84$\times$ 842 × 3 × 84 × 84
Square 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 2×3×84×842$\times$ 3$\times$ 84$\times$ 842 × 3 × 84 × 84
Tool Hang 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 2×3×240×2402$\times$ 3$\times$ 240$\times$ 2402 × 3 × 240 × 240
Transport 0.1 0.1 200 12-DOF Vel. + 2-DOF Grasping 0.05 s 4 4×3×84×844$\times$ 3$\times$ 84$\times$ 844 × 3 × 84 × 84
Push-T 0.05 [12.5,12.2] Pixels 90 2-DOF Position 0.1 s 1 1×3×96×961$\times$ 3$\times$ 96$\times$ 961 × 3 × 96 × 96
Pick Place Wall 0.6 [0.53, 0.92, 3.75] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Shelf Place 0.6 [2.32, 1.47, 4.13] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Disassemble 0.6 [0.65, 1.10, 0.61] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Stick Push 0.6 [1.03, 0.15, 0.68] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Stick Pull 0.6 [4.46, 1.44, 2.49] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Arrange Desk 0.1 [0.070, 0.071, 0.102, 0.030, 0.056, 0.148, 0.078, 0.089, 0.105, 0.039, 0.069, 0.120] rad 100 12-DOF Angle + 2-DOF Grasping 0.25 s 4 3×3×240×3203$\times$ 3$\times$ 240$\times$ 3203 × 3 × 240 × 320
Move the Stack 0.05 [0.049, 0.051, 0.057, 0.045, 0.038, 0.069] rad 100 6-DOF Angle + 1-DOF Grasping $\frac{1}{6}$ s 4 2×3×240×3202$\times$ 3$\times$ 240$\times$ 3202 × 3 × 240 × 320
@@
TABLE IX: Summary of hyperparameters for trajectory optimization and dataset. $\alpha$ is the smoothing weight, $\frac{1}{\Delta t}$ represents the frequency of policy inference, and “Demos” indicates the number of demonstration episodes. “High Dim. Obs” refers to the dimensionality of high-dimensional observations. For instance, 2×3×84×842$\times$ 3$\times$ 84$\times$ 842 × 3 × 84 × 84 corresponds to RGB observations from two cameras, each with a shape of 3×84×843$\times$ 84$\times$ 843 × 84 × 84, while 3×5123$\times$ 5123 × 512 represents point cloud data with a shape of 3×5123$\times$ 5123 × 512. “Bound” refers to $d_{\text{max}}\Delta t$ in equation (7). “Unnormalized Bound” refers to constraint bounds unnormalized to dataset scale.
[…]
Figure 13: Overlay of all 15 DiffOG trial layouts.
For real-world tasks, we randomized the initial object placements within a spatial range for each trial. An overlay visualization of all 15 DiffOG trials is shown in Fig. 13.
VI-D Evaluating DiffOG under Multiple Motion Constraints
Constraint Type No Constraint Acc. Acc. + Vel.
Base Policy Clipping Penalty-Based DiffOG DiffOG
Success Rate 0.91 0.78 0.76 0.87 0.87
Max Velocity 0.72 0.63 0.66 0.65 0.39
Max Acceleration 0.38 0.09 0.09 0.09 0.09
TABLE X: Performance on the square task under different constraint settings. DiffOG with acceleration and velocity constraints outperforms baselines that apply only acceleration constraints, highlighting its flexibility in handling multiple motion constraints.
To further demonstrate the flexibility of DiffOG in handling diverse motion-related constraints, we conducted an experiment on the square task involving multiple motion constraints, as shown in Table X. In this new experiment, we additionally incorporated an explicit velocity constraint (with a normalized bound of 0.5) into the optimization problem (6), thereby enforcing both velocity and acceleration limits during training and inference. As shown in Table X, while the addition of a velocity constraint makes the feasible solution space strictly smaller and increases the difficulty of the optimization problem, DiffOG still achieves higher success rates than baseline methods that are only subject to the acceleration constraint. This highlights DiffOG’s flexibility in integrating multiple motion constraints and its effectiveness in producing trajectories that are both well-constrained and task-successful.
VI-E Hyperparameters and Dataset
@@
The summary of hyperparameters for trajectory optimization and dataset as shown in Table IX. In the experiments of this paper, we consistently adopt symmetric constraints which is $d_{\text{max}}=-d_{\text{min}}$. Since the actions require normalization to [-1,1] during training [5], all constraints are applied on the normalized action scale for DiffOG. For simplicity, when applying constraints on the normalized action scale, the same value is used for each action dimension. Therefore, the “Bound” column contains only a single value for each task. Since the distribution of action values differs across task datasets, the bounds for each action dimension may vary after unnormalization. For Robomimic tasks [13], the action values in the dataset are already normalized to the range of [1,1][-1,1][ - 1 , 1 ]. Therefore, the bound and the unnormalized bound are identical.
@@
Since constraint clipping and penalty-based trajectory optimization do not require training, constraints can be directly applied to unnormalized actions. These constraints are equivalent to those imposed by DiffOG but are applied directly as unnormalized constraints. The parameters for penalty-based trajectory optimization are listed in Table XI. As mentioned earlier, the actions provided by the Robomimic dataset are already normalized to the range of [1,1][-1,1][ - 1 , 1 ]. The constraint bound we impose in this case is 0.10.10.1. Based on this, the violation tolerance for penalty-based optimization is set to a relatively small value of 0.0040.0040.004. For other tasks, the constraints are applied to raw action values, and the violation tolerances are determined based on the scale of unnormalized bounds provided in Table IX.
Finally, we provide the hyperparameters of the transformer used in DiffOG, as shown in Table XII.
@@
TABLE XI: Parameters of penalty-based trajectory optimization.
@@
TABLE XII: Hyperparameters of transformer encoder.
VI-F Computation of Trajectory Metrics